A 6.6 V battery is connected to a 3.3 H inductor through a 6520 ? resistor. How long will it take after the battery is connected for the energy stored in the inductor to reach 6.59E-7 J?
Answer #1
Given
voltage across battery V = 6.6 V
inductant of the inductor   L = 3.3 H
resistance of the resistor   R = 6520 Ω
energy stored in the inductor is   U = 6.59 *10 -7   J
Energy stored in the inductor is
                 E = ( 1/2) L I 2   
                2 E / L = I 2  
               2 ( 6.59 *10 -7 / 3.3 ) =    [( V / R ) ( 1 - e -Rt / L   ) ]2
                   3.9939 *10 -7   = [( 6.6 / 6520 ) ( 1 - e -Rt / L   )]2
                  0.3897 = ( 1 - e Rt / L   )2
                 ( 1 - e -Rt / L   ) = 0.6242
                       e -Rt / L =   1 - 0.6242
                                   = 0.3757
                      - Rt / L = -0.9788
                           t = 0.9788 * 3.3 / 6520
                               = 4.95 *10 -4 s